OpenStudy (anonymous):

1)Find the constant c so that a random variable X is defined by taking on values 1 , 2 , … , n , … with probabilities c*(1/4)^n. 2) Find the constant c so that a random variable X is defined assigning any interval [a,b] probability c/π⋅∫ba (1/x^2+1)) dx

OpenStudy (anonymous):

for part one i guess is 1) 1-(1/4)= 3/4 (1/4)(3/4) = 1/3 c*1/3=1 =>c=3

OpenStudy (anonymous):

for part one I don't know for sure how I'm support to do, my guess is let 3t=x 3dt=dx

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} \frac{ 1 }{ x^2+ 1}\]

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} \frac{ 1 }{ 9t^2+ 1} * 3 dt\]